OpenStudy (anonymous):

x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (uri):

Solve the bracket first −2(x + 3) = −2(x + 1) − 4 -2x-6=-2x-2=4 add 6 both sides -2x-6=-2x-2=4 +6 +6 ============ -2x=-2x+4=4 -2x+2x=4-4 They both well be cancelled.

OpenStudy (uri):

will*

OpenStudy (anonymous):

so there is no solution ?

OpenStudy (uri):

Wait let me check..

OpenStudy (uri):

−2(x + 3) = −2(x + 1) − 4 -2x-6=-2x-2-4 add 6 both sides -2x-6=-2x-2-4 +6 +6 ============ -2x=-2x -6+6 will be cancelled -2x+2x they will aslo be cancelled. No solution.

OpenStudy (anonymous):

Thanks

OpenStudy (uri):

I had put = sign by mistake except subraction sign but still no solution.